Use a calculation to estimate the mass of coal that must be burned to charge a 3.7 V 1500mAh iPhone battery assuming the electricity is generated from a coal-fired power plant, assuming the efficiency of the power plant converting heat to electricity is 0.4, the efficiency of electrical transmission from the plant to your home is 0.5, the efficiency of the phone charger at converting AC back to DC to charge your iPhone is 0.6, and assuming coal is 100% carbon.
Answer #1

von2 Twus/ Neri, carton be χ EJgicten ut 여 conventing heat into alochuerto Ne, hest tuons Ac back cion 呱.us, ク0,12 χ

Now,

0.12*x = 2 * 104

=> x = (2 * 104) / 0.12

          = 166.66 kJ

Now,

Heat of combustion of carbon is 32.8 kJ/g

Thus,

mass of carbon required, m = 166.66 / 32.8

                                              = 5.1 g